Please show me how to find the answer:

You are given two unknown point charges, Q_1 and Q
Answer #1

Since net electric field is zero,

We can say that, field produced by each charge is equal and opposite.

So, k Q1/(L/3)^2   = k Q2/(2L/3)^2

   Q1/Q2   = (L/3)^2 / (2L/3)^2  

               = 1/4